### v1

# WolBanking77: Wolof Banking Speech Intent Classification Dataset

Abstract
Intent classification models have made a lot of progress in recent years. However, previous studies primarily focus on high-resource languages datasets, which results in a gap for low-resource languages and for regions with a high rate of illiterate people where languages are more spoken than read or written. This is the case in Senegal, for example, where Wolof is spoken by around 90% of the population, with an illiteracy rate of 42% for the country. Wolof is actually spoken by more than 10 million people in West African region. To tackle such limitations, we release a Wolof Intent Classification Dataset (WolBanking77), for academic research in intent classification. WolBanking77 currently contains 9,791 text sentences in the banking domain and more than 4 hours of spoken sentences. Experiments on various baselines are conducted in this work, including text and voice state-of-the-art models. The results are very promising on this current dataset. This paper also provides detailed analyses of the contents of the data. We report baseline f1-score and word error rate metrics respectively on NLP and ASR models trained on WolBanking77 dataset and also comparisons between models. We plan to share and conduct dataset maintenance, updates and to release open-source code.
1 Introduction
Wolof is spoken in Senegal, Gambia and Mauritania by more than 10 million people. Most of the population in Senegal speaks Wolof (around 90%)  [^1] https://www.axl.cefan.ulaval.ca/afrique/senegal.htm which is essentially a language of communication between several ethnic groups. More details on Wolof language are presented in the appendix section D.
The lack of digital resources on Wolof motivated us to build the WolBanking dataset. Wolof, being an oral language, it is essential to establish voice assistants that allow the population access to digital services and to meet several challenges such as improving financial inclusion and access to digital public services. The trade sector, for instance, which plays an important role in the economy of African countries, is predominantly occupied by the informal sector [Ouattara et al., 2025]. We can also cite the agriculture, livestock, fishing, and transport sectors, which are all related to commercial activity and provide more jobs compared to the formal sector [Martínez and Short, 2022]. Recent advances in the field of artificial intelligence provide an excellent opportunity for these populations to benefit from the technology. They can gain better control over financial transactions and minimize the risk of fraud due to language barriers. In fact, the language barrier often forces business owners in the informal sector to trust a third party to consult their balance or make payments on their behalf. This potentially exposes them to the risk of fraud [Anthony et al., 2024, Ouattara et al., 2025].
However, to build such a virtual assistant, it is necessary to understand the human request to be able to provide an appropriate response. In Natural Language Processing (NLP), the field that deals with the understanding of human language, is Natural Language Understanding (NLU). More particularly in our case, we propose to work with both text and speech given that our objective is to work with African languages which are based on oral tradition. Before doing an NLU task, the first step is to understand the information contained in speech using Spoken Language Understanding (SLU). According to the World Bank Bank [2022], the rate of adult illiterate population in Senegal is 42%, which means that to facilitate access to a virtual assistant for these populations, it is essential to offer voice processing solutions. To determine the intent of a user request, Coucke et al. [2018] approach the problem as an intent detection classification task, carried out from a text or an audio transcription.
Models based on neural networks has emerged in recent years in the field of intent detection [Gerz et al., 2021, Krishnan et al., 2021, Si et al., 2023b]. However, very few datasets in African languages Conneau et al. [2022a], Mastel et al. [2023], Moghe et al. [2023], Mwongela et al. [2023], Kasule et al. [2024] have been explored. Even less in the field of e-banking for Sub-Saharan languages. Most datasets are in English Coucke et al. [2018] because development costs are very high, and substantial effort is required for their creation. This paper presents an intent classification dataset in Wolof, an African and Sub-Saharan language, based on the Banking77 dataset Casanueva et al. [2020a] which originally comprises both 13,083 customer service queries labeled with 77 intents and an audio dataset based on MINDS-14 Gerz et al. [2021] with 14 intents derived from a commercial e-banking system.
2 Related work
2.1 Existing datasets
Several datasets, both monolingual and multilingual, for NLU conversational question-answering system, have been published in the last decade. For instance, The Chatbot Corpus dataset composed of questions and answers and also The StackExchange Corpus based on the StackExchange platform both available on GitHub  [^2] https://github.com/sebischair/NLU-Evaluation-Corpora and evaluated by [Braun et al., 2017]. More recently, MINDS-14 Gerz et al. [2021] a multilingual dataset in the field of e-banking has been made publically available. When considering multilingual intent classification benchmarking, we can mention for instance the MultiATIS++ dataset Xu et al. [2020], which pertains to the aviation field, and has been expanded upon the original English ATIS dataset Price [1990] to six additional languages. More recently, the XTREME-S benchmark Conneau et al. [2022a] aims at evaluating several tasks such as speech recognition, translation, classification and retrieval. XTREME-S brings together several datasets covering 102 various languages, including 20 languages of Sub-Saharan Africa.
Concerning the Wolof language, datasets have been published in the literature, particularly in the field of NLP and Automatic Speech Recognition (ASR). We can cite some datasets that contain Wolof texts such as afriqa dataset Ogundepo et al. [2023] which deals with the question answering (QA) task, masakhaner versions 1 and 2 Adelani et al. [2021, 2022b] which targets the Name Entity Recognition (NER) task, UD_Wolof-WTB  [^3] https://github.com/UniversalDependencies/UD_Wolof-WTB or even MasakhaPOS  [^4] https://github.com/masakhane-io/lacuna_pos_ner which deals with the part-of-speech tagging (POS) task. In addition, several datasets for the machine translation (MT) task such as OPUS,  [^5] https://opus.nlpl.eu/ FLORES 200 team et al. [2022], NTREX-128 Federmann et al. [2022] and MAFAND-MT [Adelani et al., 2022a]. Datasets in the field of ASR containing Wolof such as ALFFA Gauthier et al. [2016], fleurs Conneau et al. [2022b] and more recently KALLAAMA Gauthier et al. [2024] have been published. The cited datasets cover various fairly general areas from news or religion. However, to date, there is only one text dataset dedicated to the banking sector named INJONGO Yu et al. [2025] (with other domains such as home, kitchen and dining, travel and utility). This dataset includes slot-filling and intent classification for 16 African languages including Wolof.
The authors Casanueva et al. [2020b] explored few-shot learning scenario in addition to the creation of the Banking77 dataset with 77 intents and 13,083 examples. ArBanking77 Jarrar et al. [2023] is the Arabic language version of the Banking77 dataset Casanueva et al. [2020a] in which the authors carried out simulations in low-resource settings scenario by training their model on a subset of the dataset. Other recent contributions to low-resource languages have been made, the authors Mastel et al. [2023] used Google Cloud Translation API to translate the ATIS dataset Price [1990] in Kinyarwanda and Swahili which are languages spoken by approximately 100 million people in East Africa. Moghe et al. [2023] publish a dataset called MULTI3NLU++ for several languages including Amharic. Kasule et al. [2024] propose a voice command dataset containing 20 intents in Luganda with the aim of being deployed on an IoT device.
2.2 Summary of contributions
In this work, 9,791 customer service queries from the Banking77 dataset Casanueva et al. [2020a] have been translated to French and Wolof by a team of linguistic experts from the Centre de Linguistique Appliquée de Dakar (CLAD). Additionally, this work releases another dataset based on MINDS-14 Gerz et al. [2021], with each query accompanied by audio data recorded from several speakers with varying accents and ages. The dataset covers 10 associated intentions with text and audio examples. The open source tool Lig-Aikuma Blachon et al. [2016] has been used to make the voice recordings. Our contributions include:
  • An audio dataset containing 263 sentences covering 10 intents in the banking and transport domains with varied voices and accents making this dataset the first of its kind in the Wolof language at the time of writing.
  • A text dataset with 9,791 sentences translated from English to French and Wolof covering 77 intents in banking domain.
  • Training and evaluation codes to facilitate the reproducibility of experimental benchmarks.
  • The results obtained with state-of-the-art models in various tasks such as Automatic Speech Recognition (ASR) and Intent Detection.
  • A dataset documention (datasheets) for WolBanking77.
All datasets and source codes are released under a CC BY 4.0 license to stimulate research in the field of NLP for low-resource African languages.
3 Data collection
We present in this section the main information on the creation of WolBanking77, a more detailed description can be found in Appendix E.
3.1 Textual data
[…]
Figure 1: A query translated in French and Wolof
Duplicated sentences translated in Wolof has been removed for the Intent Detection task to avoid many-to-one translations from English to Wolof. Two versions of the dataset are reported on table 2. The first version is comprised of 5k samples which is a sub-sample of the second version of 9,791 samples. Train set is 80% of both versions and test set represents 20%. Note that intents are unbalanced, the most represented intent has a frequency of 200 while the least represented one has a frequency of 24.
Some statistics were collected on the data translated into French and Wolof. Statistics on table 2 show a slightly higher number of words per query for French (83) compared to Wolof (81).
Table 1: WolBanking77 dataset statistics
WOLOF FRENCH
Min Word Count 2 2
Max Word Count 81 83
Mean Word Count 12.22 12.47
Median Word Count 10 10
Table 2: WolBanking77 dataset split
SPLIT 5k sample all
TRAIN SET 4,000 7,832
TEST SET 1,000 1,959
3.2 Audio recordings and transcriptions
The audio corpus based on MINDS-14 Gerz et al. [2021] is made up of 186 queries and 77 responses initially translated from French into English thanks to a team of linguistic experts from the Centre de Linguistique Appliquée de Dakar (CLAD). The dataset also includes a phonetic transcription of the text in Wolof in order to facilitate the correct pronunciation of the sentences by the different speakers. Additional intents have been added to the initial MINDS-14 dataset which are OPEN_ACCOUNT for information when opening a bank account, BUS_RESERVATION for reserving a transport bus, TECHNICAL_VISIT for reserving an appointment for a technical inspection of a vehicle. TRANSFER_MONEY for the intention to transfer money and AMOUNT to specify the amount to transfer. See table 3.2 for the complete list of intents of the audio dataset.
Figure 2 shows the top 5 most frequent words in the datastet after excluding stopwords. The dataset has 272 unique words and 3204 audio clips. The audio clips are in wav format using a single channel with a sampling rate of 16 kHz. Sentences have an average duration of 4815 ms. The intents ’CASH_DEPOSIT’, ’FREEZE’, ’LATEST_TRANSACTIONS’ contain the longest queries, while ’AMOUNT’, ’BUS_RESERVATION’, ’OPEN_ACCOUNT’ and ’TRANSFER_MONEY’ generally have a low query duration. The total duration of the dataset is around 4 hours and 17 minutes. Table 3: List of intents. intent domain BALANCE e-banking CASH_DEPOSIT FREEZE LATEST_TRANSACTIONS PAY_BILL OPEN_ACCOUNT TRANSFER_MONEY AMOUNT BUS_RESERVATION transport TECHNICAL_VISIT kontxaalisbëggoonajëflantexam65653737343428282727 Figure 2: Top 5 most frequent words: kont (account), xaalis (money), bëggoona (I wanted to), jëflante (operation), xam (know).
The dataset was collected thanks to the participation of students from Cheikh Anta Diop University in Dakar (UCAD), more precisely at the Faculty of Letters and Human Sciences. Each participant recorded their voice using the elicitation mode of the Lig-Aikuma software [Blachon et al., 2016]. 186 utterances were pronounced by the participants in a closed environment. The elicitation mode of the Lig-Aikuma software was installed on an Android tablet. To start a session, information about the participant are requested such as native language, region of origin (see figure 3), gender, year of birth and name. This information is subsequently stored as metadata on the tablet’s memory card. More details on demographic data and distribution information can be found in Appendix A.1.
To anonymize the name of the participants, the system generates a user_id which is used to replace the user name (more details on the ethical collection and processing of personal data are provided in Appendices A.6 and C). Then, a txt format file that contains the sentences to be pronounced is selected. One sentence at a time is presented during recording with the option to cancel recording if there is a mispronunciation at any stage. In terms of gender diversity, a total of 31 recording sessions were conducted including 14 males, 14 females and 3 unspecified. Text and audio data were split with 80% reserved for train set and 20% for test set. Before data splitting, some data preprocessing was made such as, removal of punctuation marks and numbers, lowercase text transformation. Corrupted audio files were removed from the dataset.
Dakar25.8%Diourbel12.9%Kaolack25.8%Louga6.5%Matam19.4%Thies9.7%
Figure 3: Senegal region repartition by speaker
4 Tasks & Settings
4.1 Automatic Speech Recognition
Automatic Speech Recognition (ASR) refers to the technology that enables a model to recognize and convert spoken language into text. ASR systems are widely used in various applications such as voice assistants, transcription services, customer service automation and more, whose researchers have done significant work such as Listen Attend and Spell by Chan et al. [2015] which is an end-to-end model for speech recognition based on a sequence-to-sequence neural network model. Graves et al. [2006] proposed Connectionist Temporal Classification (CTC) model used for training deep neural networks in speech recognition as well as other sequential problems where there is no explicit alignment information between the input and output. More recently, the NVIDIA team Żelasko et al. [2025] developed the Canary Flash model, which is one of the Canary Puvvada et al. [2024] variants and has the particularity of being multilingual and multitask. This model was able to achieve state-of-the-art results on several benchmarks including ASR.
Another state-of-the-art model published by Microsoft and named Phi-4-multimodal-instruct Microsoft et al. [2025] has recently been made available to address ASR tasks as well as in vision and text. Phi-4-multimodal shows an average word error rate score of 6.14% and occupies second place on Huggingface’s OpenASR leaderboard  [^6] https://huggingface.co/spaces/hf-audio/open_asr_leaderboard at the time of writing. To evaluate ASR models, Word Error Rate (WER) scores are reported in Section 5.4.
4.2 Intent Detection
Intent detection is an NLP task which consists of classifying a sentence in order to detect its intention. Several contributions have been made in the field, such as Gerz et al. [2021] who published MINDS-14 which is the first training and evaluation dataset for the intent detection task using spoken data. It includes 14 intents derived from e-banking domain, with spoken examples in 14 different languages. Si et al. [2023a] presented SpokenWOZ, a large-scale speech-text dataset for spoken Task-Oriented Dialogue (TOD) in 8 domains and 249 hours of audio. Casanueva et al. [2020b] published BANKING77, a dataset made of 13,083 examples across 77 intents, in banking domain.
In this article, we want to estimate the challenging potential of our dataset to better highlight its interest for the community by evaluating downstream tasks with increasingly sophisticated ML Workflows. We first consider classic machine learning algorithms such as k-nearest neighbor (KNN), support vector machine (SVM), linear regression (LR) and naive bayes (NB) as first baselines. The results are reported in Appendix A.4. We then consider slightly more sophisticated solutions, with the LASER sentence encoder Heffernan et al. [2022] pre-trained on several languages, including Wolof, for sentence encoding, followed by standard classification models such as MultiLayer Perceptron (MLP) and Convolution Neural Networks (CNN). We report results in Appendix A.5. Then, more advanced models based on BERT are evaluated in zero-shot-classification and few-shot-classification mode, and these same models are fine-tuned on the WolBanking77 dataset for comparison. Finally, several benchmarking of recent Small Language Models (SLM) such as Llama-3.2-3B-Instruct  [^7] https://huggingface.co/meta-llama/Llama-3.2-3B-Instruct and Llama-3.2-1B-Instruct  [^8] https://huggingface.co/meta-llama/Llama-3.2-1B-Instruct are reported. To evaluate intention detection models, f1, precision and recall metrics are reported in Section 5.4.
5 Dataset evaluation & Experiments
In this section, pre-trained models on different languages, including African ones, are presented for Intent Detection and for Automatic Speech Recognition tasks. The details of the hyperparameter settings for the experiments in this section can be found in Appendix A.2. The results show the interest of the WolBanking77 dataset for the community in order to improve intent detection for Wolof and provide an approach for extension to other low-ressource languages.
5.1 Intent detection models
The next sections present and detail the pre-trained models used in the intent detection task. Experiments for zero and few shot classification are also discussed.
5.1.1 Zero-shot classification with WolBanking77
Zero-shot text classification is a natural language processing task in which a model, trained on labeled data from certain classes, is capable of classifying text from entirely new, unseen classes without additional training. In this experiments WolBanking77 provides the new, unseen datatset. To simulate Zero-shot text classification, following pre-trained models are considered:
BERT base (uncased): A transformers model pre-trained on English introduced by Devlin et al. [2018], using a Masked Language Modeling (MLM) objective.
Afro-xlmr-large: Based on XLM-R-large Conneau et al. [2020] using MLM objective, this model has been published by Alabi et al. [2022] and pre-trained on 17 African languages but not Wolof, while still demonstrating in the original paper good scores for NER task on Wolof.
AfroLM_active_learning: Based on XLM-RoBERTa (XLM-R) using MLM objective Ogueji et al. [2021], this model has been published by [Dossou et al., 2022]. AfroLM_active_learning is a Self-Active Learning-based Multilingual Pretrained Language Model for 23 African Languages including Wolof.
mDeBERTa-v3-base-mnli-xnli: Based on DeBERTaV3-base He et al. [2021a, b], this model has been published by Moritz et al. [2022] and pre-trained on the CC100 multilingual dataset Conneau et al. [2020], Wenzek et al. [2020] with 100 different languages including Wolof. This model can perform Natural Language Inference (NLI) on 100 languages.
AfriTeVa V2 Base: A multilingual T5 model released by Oladipo et al. [2023] and pre-trained on WURA dataset Oladipo et al. [2023] covering 16 African Languages not including Wolof.
F1-score metrics for Zero-shot text classification are reported in Section 5.4.
5.1.2 Few-shot classification with WolBanking77
Given the results obtained on zero-shot classification, it seems relevant to use manually annotated data in order to improve the generalization capacity of existing models on the data presented in this article, WolBanking77. All pre-trained models cited in section 5.1.1 are selected for the few-shot classification to measure the gap between large pre-trained models with and without providing few Wolof samples from WolBanking77 for fine-tuning. F1-score metrics for zero-shot text classification are reported in Section 5.4.
5.2 ASR System
State-of-the-art ASR are chosen as baselines models pre-trained on several languages with the possibility of fine-tuning them on a specific language and domain. Below is the description of the models with more details:
Canary Flash: Canary Flash Żelasko et al. [2025] has been pre-trained on several languages (English, German, French, Spanish) and on various tasks such as ASR and translation. Canary Flash is based on Canary Puvvada et al. [2024] that is an encoder-decoder model whose encoder is based on the FastConformer model Rekesh et al. [2023] and the decoder on the Transformer architecture [Vaswani et al., 2017]. Canary has the particularity of concatenating tokenizers Dhawan et al. [2023] of different languages from SentencePiece. [^9] Google Sentencepiece Tokenizer In this reported work, English, Spanish, French and Wolof languages are concatenated with Canary’s Tokenizer. These tokens are then transformed into token embedding before feeding it into the Transformer Decoder. In addition to the tokens of each language, Canary uses 1152 tokens for special tokens representation. Canary has at the time of writing this article, 3 variants such as canary-1b, canary-1b-flash, and canary-180m-flash. Canary-1b-flash version is chosen for conducting experiments because of its multilingual support. The canary-1b-flash model was trained on 85K hours of speech data, including 31K public data (FLEURS Conneau et al. [2022b], CoVOST v2 Wang et al. [2020]) and the rest on private data. The Mozilla CommonVoice 12 dataset Ardila et al. [2020] is used as validation data for each language.
Phi-4-multimodal-instruct: Phi-4-multimodal is a multimodal Small Language Model (SLM) supporting image, text and audio within a single model. Phi-4-multimodal has the ability to handle multiple modalities without interference thanks to the Mixture of LoRAs [Hu et al., 2021]. To support multilingual inputs and outputs, the tiktoken tokenizer  [^10] Tiktoken Tokenizer is used with a vocabulary size of approximately 200K. The model is based on a Transformer decoder Vaswani et al. [2017] and supports a context length of 128K based on LongRopE [Ding et al., 2024]. Regarding the Speech/Audio modality, several modules have been introduced including the audio encoder composed of 3 CNN layers and 24 Conformer blocks [Gulati et al., 2020]. To map the dimension (1024-d) of audio features to the text embedding space (3072-d), 2 MLP layers are used in the Audio Projector module. Note that LoRAA was used on all attention and MLP layers with a rank of 320. Phi-4-multimodal was trained on 2M hours of private speech-text pairs in 8 languages. A second post-training phase by Supervised Fine Tuning (SFT) of speech/audio data pairs was conducted. For the ASR task, it includes 20k hours of private data as well as 20k hours of public data covering 8 languages. The SFT data follows the following format:
Task prompt designates the task on which the model is to be fine-tuned. Phi-4-multimodal outperforms nvidia/canary-1b-flash Żelasko et al. [2025], WhisperV3 1.5B Radford et al. [2022b], SeamlessM4T-V2 2.3B Communication et al. [2023], Qwen2-audio 8B Chu et al. [2024], Gemini-2.0-Flash Team et al. [2023] and GPT-4o OpenAI et al. [2024] on the OpenASR dataset leaderboard  [^11] https://huggingface.co/spaces/hf-audio/open_asr_leaderboard with a WER score of 6.14. Distil-whisper-large-v3.5: Distil-whisper-large-v3.5 is based on Whisper introduced by Radford et al. [2022a], which underwent training through supervised learning using 680,000 hours of labeled audio data. The authors showed that models trained with this volume of data could generalize to any dataset through zero-shot learning, meaning it can adapt without requiring fine-tuning for any specific dataset. Note that the training data encompassed 97 different languages. Several versions of Whisper followed in the last years including version 3 on which the distil-whisper-large-v3.5 model was built by knowledge-distillation following the methodology described by Gandhi et al. [2023]. Distil-whisper-large-v3.5 is trained on 98k hours of diverse filtered datasets such as Common Voice Ardila et al. [2020], LibriSpeech Panayotov et al. [2015] , VoxPopuli Wang et al. [2021] , TED-LIUM Hernandez et al. [2018], People’s Speech Galvez et al. [2021], GigaSpeech Chen et al. [2021], AMI Carletta et al. [2006], and Yodas [Li et al., 2023].
5.3 Model training
All models presented in this article are based on Pytorch library [Paszke et al., 2019]. All experiments were run on Runpod  [^12] https://www.runpod.io/, and Kaggle. [^13] https://www.kaggle.com/ P100 GPU (16GB VRAM) was used to finetune BERT base (uncased), mDeBERTa-v3-base, Afro-xlmr-large and AfroLM_active_learning. RTX 4090 GPU for AfriteVa V2. For Zero-shot and few-shot classification, RTX 2000 Ada GPU (16 GB VRAM) was used. Few-shot classification is done with SetFit huggingface library [Tunstall et al., 2022]. 2- shots refers to 2 samples per intent while 8-shots represents 8 samples per intent. All pre-trained text models are hosted on huggingface platform, they are fine-tuned using transformers library Wolf et al. [2020] except for Llama-3.2 which is fine-tuned using torchtune library [torchtune maintainers and contributors, 2024].
MPL and CNN models are trained from scratch on 200 epochs with a single 800-dimension hidden layer, ReLU activation layer and a fully connected output layer for MLP. For CNN, a Conv1d followed by a ReLU activation layer and a fully connected layer for output classification. The Wolbanking77 data was structured in the form of a prompt illustrated in Appendix A.3 in order to allow the Llama-3.2 model to generate the correct intent.
Canary Flash and Phi-4-multimodal-instruct ASR models are fine-tuned on the Wolbanking77 audio dataset using respectively the NVIDIA NEMO framework Kuchaiev et al. [2019] and huggingface Transformers on A100 SXM GPU (80GB VRAM). In the other hand, Distil-whisper-large-v3.5 is fine-tuned using the Huggingface Transformers framework on RTX 2000 Ada GPU (16 GB VRAM). All ASR models are fine-tuned on 1000 steps.
5.4 Results and Discussions
Figure 4: Zero-shot, few-shot and fine-tuning (FT) results (in % for F1-score) of multilingual pre-trained models on WolBanking77.
BertBaseAfroXLMRAfroLMmDeBERTa-v3AfritevaV20505010010011110.20.23311202026261717101014144646474741413535282872727979727269695555F1-scoreScores on 5k data0-shot2-shots8-shotsFT
Figure 4 shows weighted average F1-score results for zero-shot and few-shot classification. AfroXLMR model outperforms all multilingual pre-trained models on few-shot and fine-tuning (with an F1-scores of 79%) on 5k samples followed by BertBase and AfroLM. This results also show that existing models even when pre-trained on Wolof do not perform well and that our dataset provides specific challenges.
Figure 5: Zero-shot and fine-tuning (FT) results (in %) on multilingual pre-trained models.
BertBaseAfroXLMRAfroLMmDeBERTa-v3AfritevaV2050501001002.32.3220.40.41.51.50.50.555555757565652525050F1-scoreScores on all samples0-shotFT
Experiments have been conducted on the entire WolBanking77 dataset and reported on figure 5. Similar to previous experiments conducted on 5k sub-sample of the WolBanking77 dataset, results shows that AfroXLMR achieves slightly better scores compared to BertBase and AfroLM. These results show that intent detection for low-resource languages remains a challenging task even for state-of-the-art small language models, as shown by the results reported in table LABEL:tab:slm.
Table 4: Precision, recall and f1-score scores for Small Language Models on WolBanking77. Best models are highlighted in lightgray.
5k samples All samples
Model Precision Recall F1 Precision Recall F1
Llama-3.2-1B-Instruct 0.74 0.71 0.72 0.52 0.45 0.46
[…]
A comparison is also conducted for ASR models on the audio part of WolBanking77 dataset. Special characters have been removed from the text and then reduced to lowercase. We can observe from the results with 4 hours of speech data in Table 5 that Canary-1b-flash with a wer score of 0.59% outperforms Phi-4-multimodal-instruct (wer 3.1%) and more particularly Distil-whisper-large-v3.5 (wer 4.63%). All pre-trained ASR models have been fine-tuned on the audio part of WolBanking77 dataset with 1000 steps. The results obtained following the experiments show that it is possible to obtain good results with little data, which is favorable for our WolBanking77 dataset and in low-resource language contexts.
Table 5: Word Error Rates (WER) with 4 hours of speech. Training time is reported in minutes.
Model Parameters Training time Steps WER
Phi-4-multimodal-instruct 5.6B 32 1000 3.1%
Distil-whisper-large-v3.5 756M 44 1000 4.63%
[…]
6 Long-term Support and Future Work
Long term support includes updates to the text and audio recordings, improvements of the simulation setup script, and possible bug fixes. Free access is offered to various versions of the dataset and source code under a CC BY 4.0 license to accelerate research in the field. It is planned to add more audio recordings in various environments to make the models more robust. Text data of possible responses to each intent will be shared for possible usage in Text-To-Speech. The dataset can also be annotated for slot-filling task.
7 Conclusion
Access to financial services or public transportation services can be facilitated for illiterate people by providing them with voice interfaces in their language of communication. In this paper, an intent detection dataset is presented in Wolof language in two modalities which are voice and text. Additionally, dataset description and benchmarks are presented. WolBanking77 is published and shared freely under CC BY 4.0 license along with the code and datasheets Gebru et al. [2021] with the aim of inspiring low budget research into low-resource languages.
Acknowledgments and Disclosure of Funding
This work is supported by the Partnership for Skills in Applied Sciences, Engineering and Technology (PASET) - Regional Scholarship and Innovation Fund (RSIF) under Grant No.:B8501L10014. We also thank the entire CLAD team and in particular: Professor Souleymane Faye and all the translators who participated in this project.
Appendix A Technical Appendices and Supplementary Material
A.1 WolBanking77 supplementary details
Does the partition illustrated in figure 3 correspond to the population of each region?
The score presented is not representative of the population of each region. We tried to cover various ethnic groups in the country in order to have diverse accents.
Are these regions fully representative of the linguistic variation that occurs in the Wolof language across Senegal?
Wolof has several dialects, including Bawol, Kajoor, Jolof, and Jander; This dialectal diversity does not prevent inter-understanding. However, there is a lesser degree of inter-understanding between these varieties and that spoken by the Lebous (an ethnic group living mainly in the Cape Verde peninsula, i.e. in certain localities in the Dakar region).
Is the Wolof language as used in Mauritania and Gambia substantially distinct from the Senegalese variety, such that the demonstrated performance on the audio dataset may not generalise to these Wolof-speaking populations?
The Wolof language used in Mauritania and Gambia is not substantially different. However, the differences are to be seen rather in linguistic evolutions and in particular with the borrowings from foreign languages present, in particular Arabic and English. These differences may pose a problem for the generalization of models trained on the audio data that do not yet cover these variants of Wolof. This is also present in the accent of the Wolof locutors which highly differs under the effect of the official language (English or Arabic). This has a strong impact on Wolof words pronunciation.
It would be useful to know the age distribution of the speakers.
The age distribution of the participants are presented on table 6.
It would also be useful to know the intersection of age and literacy in the population. If older speakers, for example, were proportionally less literate in the language, this might argue for strengthening the representation of older voices in the dataset, given that this development is aimed at improving accessibility through voice-interactive technology.
According to statistics (from Agence Nationale de la Statistique et de la Démographie ANSD [2013]), the literacy rate is higher among the younger populations of 10-14 years old and 15-19 years old, with 58.1% and 64.1% respectively. However, the literacy rate decreases among older populations over 70 years of age (between 13% and 20%). In view of these figures, we intend to take the representation of the older population in future versions of the dataset.
Table 6: Age distribution of the speakers
Stat Age
mean 26
std 4
min 22
25% 23
50% 25
75% 27
max 36
Is the digital infrastructure (mobile or desktop internet services) sufficient in all regions to support the deployment of the proposed future technology?
The mobile phone and internet coverage is now very high in Senegal, the WhatsApp application for example is used throughout the country.
A.2 Hyperparameters
Table 7: Hyperparameters for NLP models. mDeBERTa-v3* and AfritevaV2 have same hyperparameters.
Model bert-base-uncased AfroXLMR mDeBERTa-v3* AfroLM Llama3.2
Learning Rate 2e-05 2e-05 2e-05 2e-05 3e-4
Train Batch Size 32 4 8 16 4
Eval Batch Size 32 8 8 8 4
Warmup ratio 0.1 0.1 0.1 0.1 -
# epochs 20 20 20 20 20
NLP models hyperparameters are reported on table 7, AdamW_torch_fused is the default optimizer for NLP models except for Llama3.2 that is AdamW.
Table 8: Hyperparameters for ASR models.
Model Canary-1b-flash Distil-whisper-large-v3.5 Phi-4
Learning Rate 3e-4 1e-5 1e-4
Train Batch Size - 8 16
Eval Batch Size 8 8 -
Gradient Accumulation Steps - 1 1
Lr Scheduler Warmup Steps 2500 500 -
Optimizer AdamW AdamW AdamW_torch
# steps 1000 1000 1000
Table 8 shows the hyperparameters used to train Phi-4-multimodal-instruct, Distil-whisper-large-v3.5 and Canary-1b-flash. Hyperparameters has been chosen by following HuggingFace Wolf et al. [2020] and NVIDIA Nemo Kuchaiev et al. [2019] documentations. AdamW has been used as a default optimizer.
Table 9: Best hyperparameters for LASER3+CNN.
Model LASER3+CNN
Learning Rate 0.013364046097018405
Last size of non classification layer 128
Batch Size 2
A.3 Text prompt for Llama3.2
This prompt format was used to train the Llama3.2 model.
{text} is a variable that can be one of the following sentences for example: Dama wara yónnee xaalis, ndax mën naa jëfandikoo sama kàrtu kredi? (I need to transfer some money, can I use my credit card?) Jotuma xaalis bi ma ñu ma waroon a delloo. (I have not received a refund.) Ci lan ngeeni jëfandikoo kàrt yi ñuy sànni? (What do you use disposable cards on?) Ndax amna anam buma mëna toppe lii? (Was there a way for me to get tracking for that?) Ban diir la wara am ngir yonnee xaalis Etats-Uni? (How long does it take for deliver to the US?)
A.4 Reference scores from classic Machine Learning Techniques
Machine learning models such as KNN, SVM, Logistic Regression (LR) and Naive Bayes (NB) with Bag-Of-Words are used as baselines as well as CNN and MPL with LASER3 as sentence encoder. Table LABEL:tab:ml shows ML baselines comparison. Results shows that Bag of Words combined with Linear Regression outperforms the other ML models with an F1-score of 53%. However, LR and SVM achieve same results on 5k split.
Table 10: ML baselines models precision, recall and f1-score results on WolBanking77. Best models are highlighted in lightgray.
5k samples All samples
Model Precision Recall F1 Precision Recall F1
BoW+KNN 0.48 0.39 0.39 0.37 0.31 0.31
BoW+SVM 0.70 0.69 0.68 0.49 0.48 0.48
[…]
BoW+NB 0.54 0.51 0.50 0.28 0.26 0.26
A.5 Pretrained Sentence Encoder
LASER3: Heffernan et al. [2022] propose an improvement of LASER by training the model on several languages with the aim of encoding them in the same representation space. The method used consists of training a teacher-student model by combining supervised and self-supervised approaches with the aim of training the model on low-resource languages. This allows to cover 50 African languages including Wolof. Teacher-student training is used in order to avoid training a new model from scratch each time we need to encode a new language. Some of the African languages come from the Masakhane project  [^14] https://www.masakhane.io/ as well as the EMNLP’22 workshop. [^15] https://www.statmt.org/wmt22/large-scale-multilingual-translation-task.html The authors made some changes to the LASER architecture such as replacing BPE with SPM (SentencePiece Model), which is an unsupervised text tokenizer and detokenizer that does not depend on language-specific pre or postprocessing, as well as adding an upsampling step for low-resource languages. The LASER sentence encoder trained on the public corpus OPUS  [^16] https://opus.nlpl.eu/ is used as a teacher model and renamed by LASER2, then the student model is named LASER3. A student model is trained for each language covered in Masked Language Modeling mode, then a multilingual distillation approach is applied by optimizing the cosine loss between the embeddings generated by the teacher and the student. The student architecture has been replaced by a 12-layer transformer instead of a 6-layer BiLSTM of the original LASER. Concerning Wolof language, the model was trained on 21k bitexts which are pairs of texts in two different languages that are translations of each other and 94k sentences using training distillation in addition to Masked Language Modeling. The results of the experiments showed a clear improvement of LASER3 encoder on Wolof with a score of 6.03 xsim error rate (a margin-based similarity score by Artetxe and Schwenk [2019]) compared to the original LASER encoder which produced a score of 70.65.
Pre-trained sentence encoder LASER3 that vectorize each sentence and all possible combinations of classification models (MPL and CNN) are compared in the reported results on Table LABEL:tab:laser_sent_enc. Note that all classification models have been trained from scratch. For all the evaluations, punctuation are removed and each sentence transformed in lowercase.
Table 11: LASER3 with classification models precision, recall and f1-score results on WolBanking77. LASER3+CNN (tuning) denotes LASER3+CNN with CNN hyperparameter tuning. Best models are highlighted in lightgray.
5k samples All samples
Model Precision Recall F1 Precision Recall F1
[…]
LASER3+CNN 0.01 0.05 0.01 0.01 0.05 0.01
LASER3+CNN (tuning) 0.53 0.50 0.49 0.33 0.32 0.32
Table LABEL:tab:laser_sent_enc shows weighted average results of LASER3 with classification models (MLP, CNN). LASER+MLP achieves best results with F1-score of 55% on 5k samples and 42% on all samples compared to LASER+CNN which produced poor results with a low F1-score of 1%. To improve the LASER+CNN model, we have tuned the CNN model hyperparameters relying on the Ray Tune library [Liaw et al., 2018]. The tuned hyperparameters are: the size of the last internal representation before the classification layer, the learning rate (lr) and the batch size. We have run the tuning on 10 different trials, and for each trial, Ray Library randomly sampled a combination of parameters using ASHAScheduler which terminates early the trials performing badly. Best hyperparameters obtained are reported on table 9. This leads to the results reported on table LABEL:tab:laser_sent_enc.
A.6 Consent form
Figure 6 presents the consent form (in French) presented to participants before the recording sessions.
Figure 6: Consent form
Appendix B Limitations
We can note some limitations in our dataset such as intention classes imbalance which can push the model to favor majority classes to the detriment of minority ones. During the ASR evaluation, we noticed spelling errors on rare words which can be corrected with a language model.
Appendix C Ethics Statement
To protect the anonymity of participants, Personal data such as speaker name, location and androidID have been removed in the dataset. We notified each participant that the audio collected will be used as a public dataset for research purposes. Each participant read and signed a consent form in which they declare their free and informed acceptance to participate as a subject in the study as part of the data collection proposed in this project as described in section A.6. We paid the amount of \$850 to the linguists for the translation and phonetic transcription of all the sentences in Wolof.
Appendix D Wolof language
Wolof is a member of the West Atlantic sub-branch of the Niger-Congo language family. A small minority from various other ethnic groups have adopted Wolof as their first language, while more than half of non-native speakers use Wolof as their second or third language. Together, these three groups of Wolof speakers represent about 90% of the population, making Senegal one of the most linguistically unified nations in West Africa [Mbodj, 2014]. Wolof is also spoken by the majority of the population in Gambia [Ndione, 2013]. Genetically related languages are Pulaar and Serere, which are all languages of the Niger-Congo phylum, and of the Atlantic branch Ndione [2013] see figure 7.
Figure 7: The main languages of Senegal Leclerc [2023]
To construct sentences in Wolof, there are several processes such as 8 class markers for words in singular form and 2 class markers for words in plural form. These markers are positioned in front of or right after the word such as: nit ki (to designate a person), nit ñi (to designate several persons). In addition, we also find other construction processes such as consonantal alternation, pre-nasalization and suffixation. For some derivations, there is the simultaneous presence of several of these processes. For example, pre-nasalization can be accompanied by suffixation [Ndione, 2013].
The conjugation in Wolof is achieved thanks to an invariable lexical base to which are added affixes holding the markers of IPAM (Index; Person Aspect-time; Mode) [Perrin, 2005, Ndione, 2013]. These inflectional markers highlight both semantic roles and grammatical relations. We mainly distinguish two types of verbs in Wolof, action verbs (expressing an action or a fact carried out or suffered by the subject) and state verbs (giving characteristics to elements, they describe a state, a way of being). There are ten conjugations in Wolof Perrin [2005], Ndione [2013] such as the perfect, the aorist, the presentative, the emphatic of the verb, the emphatic of the subject, the emphatic of the complement, the negative, the emphatic negative, the imperative and the obligative.
Depending on the grammar, in Wolof, there is generally only one auxiliary, "di", with its variants "d" and "y" which mark the unaccomplished. The variant "d" appears with the time and the negation markers (doon, daa, daan, dee, du).
Appendix E Datasheets for WolBanking77
E.1 Motivation
For what purpose was the dataset created? Was there a specific task in mind? Was there a specific gap that needed to be filled? Please provide a description.
This dataset was created with the aim of propelling research on intent detection in Wolof which is a language spoken in West Africa in order to allow illiterate people to be able to interact with digital systems through the voice channel. We introduce a text dataset and audio dataset including utterances accompanied by their transcription with their corresponding intent in order to be able to simulate a dialogue between the user and the system.
Who created the dataset (e.g., which team, research group) and on behalf of which entity (e.g., company, institution, organization)?
This dataset is created by researchers at Université Cheikh Anta Diop of Dakar and Université Gaston Berger of Saint-Louis in Senegal.
Who funded the creation of the dataset?
Funding was provided by the Partnership for skills in Applied Sciences, Engineering and Technology (PASET) and Regional Scholarship and Innovation Fund (RSIF).
E.2 Composition
What do the instances that comprise the dataset represent (e.g., documents, photos, people, countries)? Are there multiple types of instances (e.g., movies, users, and ratings; people and interactions between them; nodes and edges)? Please provide a description.
The dataset includes text and audio data for intent classification. For every sentence, text and annotations are provided.
How many instances are there in total (of each type, if appropriate)?
For the audio version, there are 263 instances covering 10 intents in total. 4 hours and 17 minutes of audios from spoken sentences. For the text version, there are 9,791 instances covering 77 intents in total.
Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set? If the dataset is a sample, then what is the larger set? Is the sample representative of the larger set (e.g., geographic coverage)? If so, please describe how this representativeness was validated/verified. If it is not representative of the larger set, please describe why not (e.g., to cover a more diverse range of instances, because instances were withheld or unavailable).
The dataset contain all possible instances.
What data does each instance consist of? “Raw” data (e.g., unprocessed text or images)or features? In either case, please provide a description.
Each instance consists of the text associated with intent label. The audio data contains transcriptions of the corresponding audio files.
Is there a label or target associated with each instance? If so, please provide a description.
Yes, there is a label intent class for each instance in the dataset.
Is any information missing from individual instances? If so, please provide a description, explaining why this information is missing (e.g., because it was unavailable). This does not include intentionally removed information, but might include, e.g., redacted text.
Everything is included. There is no missing data.
Are relationships between individual instances made explicit (e.g., users’ movie ratings, social network links)? If so, please describe how these relationships are made explicit.
Each instance in WolBanking77 is relatively independent.
Are there recommended data splits (e.g., training, development/validation,testing)? If so, please provide a description of these splits, explaining the rationale behind them.
There are 80% of the data for training and 20% for validation.
Are there any errors, sources of noise, or redundancies in the dataset? If so, please provide a description.
The audio recordings were made in an uncontrolled environment and therefore contain background noise.
Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)?
WolBanking77 is self-contained.
Does the dataset contain data that might be considered confidential (e.g., data that is protected by legal privilege or by doctor-patient confidentiality, data that includes the content of individuals’ non-public communications)?If so, please provide a description.
No.
Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety? If so, please describe why.
No.
Does the dataset identify any subpopulations (e.g., by age, gender)? If so, please describe how these subpopulations are identified and provide a description of their respective distributions within the dataset.
The dataset identify subpopulations by age, gender and origin. The subpopulations are identified when participants fill the identification form by giving there age, gender and origin before starting a new recording session.
Is it possible to identify individuals (i.e., one or more natural persons), either directly or indirectly (i.e., in combination with other data) from the dataset? If so, please describe how.
No it is not possible to identify individuals from the dataset, names are removed from the dataset.
Does the dataset contain data that might be considered sensitive in any way (e.g., data that reveals race or ethnic origins, sexual orientations, religious beliefs, political opinions or union memberships, or locations; financial or health data; biometric or genetic data; forms of government identification, such as social security numbers; criminal history)? If so, please provide a description.
The dataset contain regions of origin showing that different voices and accents have been recorded to cover as more ethnic as possible and to make models robust to different accents.
E.3 Collection Process
How was the data associated with each instance acquired? Was the data directly observable (e.g., raw text, movie ratings), reported by subjects (e.g., survey responses), or indirectly inferred/derived from other data (e.g., part-of-speech tags, model-based guesses for age or language)? If the data was reported by subjects or indirectly inferred/derived from other data, was the data validated/verified? If so, please describe how.
Dataset collection process has been reported in section 3.
What mechanisms or procedures were used to collect the data (e.g., hardware apparatuses or sensors, manual human curation, software programs, software APIs)? How were these mechanisms or procedures validated?
Data has been recorded using Lig-Aikuma Android software [Blachon et al., 2016].
If the dataset is a sample from a larger set, what was the sampling strategy (e.g., deterministic, probabilistic with specific sampling probabilities)?
Wolbanking77 is not sampled from a larger set.
Who was involved in the data collection process (e.g., students, crowdworkers, contractors) and how were they compensated (e.g., how much were crowdworkers paid)?
The data was collected with the help of students and professional translators. Data collection cost is around \$5000.
Over what timeframe was the data collected? Does this timeframe match the creation timeframe of the data associated with the instances (e.g., recent crawl of old news articles)? If not, please describe the timeframe in which the data associated with the instances was created.
Our data collection started in March 2024. The contents of our dataset are independent of the time of collection.
Were any ethical review processes conducted (e.g., by an institutional review board)? If so, please provide a description of these review processes, including the outcomes, as well as a link or other access point to any supporting documentation.
Not applicable.
Did you collect the data from the individuals in question directly, or obtain it via third parties or other sources (e.g., websites)?
The text data is a translated version of Banking77 and the audio version from MINDS14.
Were the individuals in question notified about the data collection? If so, please describe (or show with screenshots or other information) how notice was provided, and provide a link or other access point to, or otherwise reproduce, the exact language of the notification itself.
A consent form shown in section A.6 has been provided to each individuals.
Did the individuals in question consent to the collection and use of their data? If so, please describe (or show with screenshots or other information) how consent was requested and provided, and provide a link or other access point to, or otherwise reproduce, the exact language to which the individuals consented.
Yes, see previous question.
If consent was obtained, were the consenting individuals provided with a mechanism to revoke their consent in the future or for certain uses? If so, please provide a description, as well as a link or other access point to the mechanism (if appropriate).
Yes, see A.6.
Has an analysis of the potential impact of the dataset and its use on data subjects (e.g., a data protection impact analysis) been conducted? If so, please provide a description of this analysis, including the outcomes, as well as a link or other access point to any supporting documentation.
No.
E.4 Preprocessing/cleaning/labeling
Was any preprocessing/cleaning/labeling of the data done (e.g., discretization or bucketing, tokenization, part-of-speech tagging, SIFT feature extraction, removal of instances, processing of missing values)? If so, please provide a description. If not, you may skip the remaining questions in this section. Dataset construction and cleaning has been reported in section 3. Text tokenization and removal of duplicated sentences has been done for NLP task.
Was the “raw” data saved in addition to the preprocessed/cleaned/labeled data (e.g., to support unanticipated future uses)? If so, please provide a link or other access point to the “raw” data. No. The dataset does not contain all the raw data and metadata.
Is the software used to preprocess/clean/label the instances available? If so, please provide a link or other access point.
Python programming language has been used to clean the data. All the code sources will be shared.
E.5 Uses
Has the dataset been used for any tasks already? If so, please provide a description.
Not yet.
Is there a repository that links to any or all papers or systems that use the dataset? If so, please provide a link or other access point.
No yet.
What (other) tasks could the dataset be used for?
The dataset can be used for anything related to intent classification as well as to train speech and text models.
Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses? For example, is there anything that a dataset consumer might need to know to avoid uses that could result in unfair treatment of individuals or groups (e.g., stereotyping, quality of service issues) or other risks or harms (e.g., legal risks, financial harms)? If so, please provide a description. Is there anything a dataset consumer could do to mitigate these risks or harms?
Not applicable.
Are there tasks for which the dataset should not be used? If so, please provide a description.
Not applicable.
E.6 Distribution
Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created?
The dataset will be distributed on the internet and publicly available.
How will the dataset will be distributed (e.g., tarball on website, API, GitHub)? Does the dataset have a digital object identifier (DOI)?
The dataset will be distributed on tarball on a website. DOI will also be available.
When will the dataset be distributed?
As soon as paper publication.
Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)? If so, please describe this license and/or ToU, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms or ToU, as well as any fees associated with these restrictions.
WolBanking77 is distributed under the CC BY 4.0  [^17] https://creativecommons.org/licenses/by/4.0/legalcode.en license. This license requires that reusers give credit to the creator. It allows reusers to distribute, remix, adapt, and build upon the material in any medium or format, even for commercial purposes.
Have any third parties imposed IP-based or other restrictions on the data associated with the instances? If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms, as well as any fees associated with these restrictions.
No.
Do any export controls or other regulatory restrictions apply to the dataset or to individual instances? If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any supporting documentation.
No.
E.7 Maintenance
Who is supporting/hosting/maintaining the dataset?
Maintenance will be supported by the authors.
How can the owner/curator/manager of the dataset be contacted (e.g., email address)?
Authors can be contacted by their email address.
Is there an erratum? If so, please provide a link or other access point.
Any updates will be shared via Kaggle.
Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)? If so, please describe how often, by whom, and how updates will be communicated to users (e.g.,mailing list,GitHub)?
If necessary, any updates will be released via Kaggle.
If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances (e.g., were the individuals in question told that their data would be retained for a fixed period of time and then deleted)? If so, please describe these limits and explain how they will be enforced.
N/A.
Will older versions of the dataset continue to be supported/hosted/maintained? If so, please describe how. If not, please describe how its obsolescence will be communicated to dataset consumers.
Older versions of the dataset will be kept.
If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so? If so, please provide a description. Will these contributions be validated/verified? If so, please describe how. If not, why not? Is there a process for communicating/distributing these contributions to dataset consumers? If so, please provide a description.
Yes. Please contact the authors of this paper for building upon this dataset.