### v3

# WolBanking77: Wolof Banking Speech Intent Classification Dataset

Abstract
Intent classification models have made a significant progress in recent years. However, previous studies primarily focus on high-resource language datasets, which results in a gap for low-resource languages and for regions with high rates of illiteracy, where languages are more spoken than read or written. This is the case in Senegal, for example, where Wolof is spoken by around 90% of the population, while the national illiteracy rate remains at of 42%. Wolof is actually spoken by more than 10 million people in West African region. To address these limitations, we introduce the Wolof Banking Speech Intent Classification Dataset (WolBanking77), for academic research in intent classification. WolBanking77 currently contains 9,791 text sentences in the banking domain and more than 4 hours of spoken sentences. Experiments on various baselines are conducted in this work, including text and voice state-of-the-art models. The results are very promising on this current dataset. In addition, this paper presents an in-depth examination of the dataset’s contents. We report baseline F1-scores and word error rates metrics respectively on NLP and ASR models trained on WolBanking77 dataset and also comparisons between models. Dataset and code available at: wolbanking77.
1 Introduction
Wolof is spoken in Senegal, Gambia and Mauritania by more than 10 million people. Most of the population in Senegal speaks Wolof (around 90%)  [^1] https://www.axl.cefan.ulaval.ca/afrique/senegal.htm which is essentially a language of communication between several ethnic groups. More details on Wolof are presented in the appendix section 11.
The lack of digital resources for Wolof motivated us to build the WolBanking77 dataset. Wolof, being an oral language, it is essential to establish voice assistants that allow the population access to digital services and help address challenges such as enhancing financial inclusion and access to digital public services. The trade sector, for instance, which plays an important role in the economy of African countries, is predominantly driven by the informal economy [Ouattara et al., 2025]. We can also cite the agriculture, livestock, fishing, and transport sectors, which are all related to commercial activity and generate more employment compared to the formal sector [Martínez and Short, 2022]. Recent advances in artificial intelligence provide an good opportunity for these populations to benefit from digital technology. They can gain better control over financial transactions and minimize the risk of fraud due to language barriers. In fact, the language barrier often compel business owners in the informal sector to rely on a third party to consult their account balances or make payments on their behalf. This potentially exposes them to the risk of fraud [Anthony et al., 2024, Ouattara et al., 2025].
However, to build such a virtual assistant, it is necessary to understand human requests in order to provide appropriate responses. In Natural Language Processing (NLP), the field that deals with the understanding of human language, is Natural Language Understanding (NLU). In our case, we focus on both text and speech given that our objective is to work with African languages rooted in oral traditions. Before doing an NLU task, the first step is to understand the information contained in speech using Spoken Language Understanding (SLU). According to the World Bank Bank [2022], the rate of adult illiterate (see Appendix 8.4) population in Senegal is 42%, which means that to facilitate access to a virtual assistant for these populations, it is essential to offer voice processing solutions. To determine the intent of a user request, Coucke et al. [2018] frame the problem as an Intent Detection (ID) classification task, performed either directly from text or from audio transcriptions.
Models based on neural networks has emerged in recent years in the field of ID [Gerz et al., 2021, Krishnan et al., 2021, Si et al., 2023b]. However, very few datasets in African languages Alexis et al. [2022], Mastel et al. [2023], Moghe et al. [2023], Mwongela et al. [2023], Kasule et al. [2024] have been explored. Even less in the field of e-banking for Sub-Saharan languages. Most existing datasets are in English Coucke et al. [2018], as their development requires significant financial and human resources. In this paper, we present an intent classification dataset in Wolof, an African and Sub-Saharan language, based on the Banking77 dataset Casanueva et al. [2020] which originally consists of 13,083 customer service queries labeled with 77 intents. In addition, our work also draws on the MINDS-14 dataset Gerz et al. [2021], which contains 14 intents derived from a commercial e-banking system and includes an audio component.
2 Related work
2.1 Existing datasets
Several datasets, both monolingual and multilingual, for NLU conversational question-answering system, have been published in the last decade. For instance, The Chatbot Corpus dataset composed of questions and answers and also The StackExchange Corpus based on the StackExchange platform both available on GitHub  [^2] https://github.com/sebischair/NLU-Evaluation-Corpora and evaluated by [Braun et al., 2017]. More recently, MINDS-14 Gerz et al. [2021] a multilingual dataset in the field of e-banking has been made publicly available. When considering multilingual intent classification benchmarking, we can mention for instance the MultiATIS++ dataset Xu et al. [2020], which pertains to the aviation field, and has been expanded upon the original English ATIS dataset Price [1990] to six additional languages. More recently, the XTREME-S benchmark Alexis et al. [2022] aims at evaluating several tasks such as speech recognition, translation, classification and retrieval. XTREME-S brings together several datasets covering 102 various languages, including 20 languages of Sub-Saharan Africa.
With regard to the Wolof language, datasets have been published in the literature, particularly in the field of NLP and Automatic Speech Recognition (ASR). For example, some datasets include Wolof texts such as afriqa dataset Ogundepo et al. [2023] which deals with the question answering (QA) task, masakhaner versions 1 and 2 Adelani et al. [2021, 2022b] which targets the Name Entity Recognition (NER) task, UD_Wolof-WTB  [^3] https://github.com/UniversalDependencies/UD_Wolof-WTB or even MasakhaPOS  [^4] https://github.com/masakhane-io/lacuna_pos_ner which addresses the part-of-speech tagging (POS) task. In addition, several datasets for the machine translation (MT) task such as OPUS,  [^5] https://opus.nlpl.eu/ FLORES 200 team et al. [2022], NTREX-128 Federmann et al. [2022] and MAFAND-MT [Adelani et al., 2022a]. In the field of ASR, several datasets containing Wolof have also been published, including ALFFA Gauthier et al. [2016], fleurs Conneau et al. [2023] and more recently KALLAAMA Gauthier et al. [2024]. The cited datasets cover fairly general domains, such as news and religion. However, to date, there is only one text dataset dedicated to the banking sector named INJONGO Yu et al. [2025] (with other domains such as home, kitchen and dining, travel and utility). This dataset includes slot-filling and intent classification tasks for 16 African languages including Wolof.
The authors Casanueva et al. [2020] explored few-shot learning scenario in addition to introducing the Banking77 dataset, which contains 77 intents and 13,083 examples. ArBanking77 Jarrar et al. [2023] is the Arabic language version of the Banking77 dataset Casanueva et al. [2020] in which the authors conducted simulations in low-resource settings scenario by training their model on a subset of the dataset. Other recent contributions to low-resource languages have been made, the authors Mastel et al. [2023] used Google Cloud Translation API to translate the ATIS dataset Price [1990] in Kinyarwanda and Swahili which are languages spoken by approximately 100 million people in East Africa. Similarly, Moghe et al. [2023] introduced the MULTI3NLU++ dataset for several languages including Amharic. Kasule et al. [2024] proposed a voice command dataset containing 20 intents in Luganda, designed for deployment on IoT devices.
2.2 Summary of contributions
In this work, 9,791 customer service queries from the Banking77 dataset Casanueva et al. [2020] was translated to French and Wolof by a team of linguistic experts from the Centre de Linguistique Appliquée de Dakar (CLAD). Additionally, this work introduces another dataset based on MINDS-14 Gerz et al. [2021], in which each query is paired with audio recordings from multiple speakers with diverse accents and ages (see Appendix 8.3). The dataset includes 10 intents represented with both text and audio examples. The open source tool Lig-Aikuma Blachon et al. [2016] was used to produce the voice recordings. Our contributions include:
  • An audio dataset with 263 sentences covering 10 intents in the banking and transport domains, including diverse voices and accents, making it the first of its kind in the Wolof language at the time of writing.
  • A text dataset with 9,791 sentences translated from English to French and Wolof covering 77 intents in banking domain.
  • The results obtained with state-of-the-art models in various tasks such as Automatic Speech Recognition (ASR) and Intent Detection (ID). We also provide training and evaluation code to support the reproducibility of experimental benchmarks.
  • A dataset documentation (datasheets) for WolBanking77.
All datasets and source codes are released under a CC BY 4.0 license to stimulate research in the field of NLP for low-resource African languages.
3 Data collection
In this section, we present the main information on the creation of WolBanking77, a more detailed description can be found in Appendix 12.
3.1 Textual data
[…]
Figure 1: A query translated in French and Wolof
Duplicated sentences translated in Wolof was removed for the ID task to avoid many-to-one translations from English to Wolof. Two versions of the dataset are reported on table 2. The first version is comprised of 5k samples which is a sub-sample of the second version of 9,791 samples. Train set is 80% of both versions and test set represents 20%. Note that intents are unbalanced, the most represented intent has a frequency of 200 while the least represented one has a frequency of 24.
Some statistics were collected on the data translated into French and Wolof. Statistics on table 2 show a slightly higher number of words per query for French (83) compared to Wolof (81).
Table 1: WolBanking77 dataset statistics
WOLOF FRENCH
Min Word Count 2 2
Max Word Count 81 83
Mean Word Count 12.22 12.47
Median Word Count 10 10
Table 2: WolBanking77 dataset split
SPLIT 5k sample all
TRAIN SET 4,000 7,832
TEST SET 1,000 1,959
3.2 Audio recordings and transcriptions
The audio corpus based on MINDS-14 Gerz et al. [2021] consists of 186 queries and 77 responses, initially translated from French into English by a team of linguistic experts from the Centre de Linguistique Appliquée de Dakar (CLAD). The dataset also includes a phonetic transcription of the Wolof text to facilitate the correct pronunciation of sentences by different speakers. Additional intents was added to the initial MINDS-14 dataset, namely: OPEN_ACCOUNT, for information related to opening a bank account, BUS_RESERVATION, for booking a transport bus (see Appendix 8.5), TECHNICAL_VISIT for scheduling a vehicle inspection appointment. TRANSFER_MONEY for the intent to transfer money and AMOUNT to specify the amount to be transferred. See table 3.2 for the complete list of intents in the audio dataset.
Figure 2 shows top 5 most frequent words in the datastet after excluding stopwords. The dataset contains 272 unique words and 3,204 audio clips. The audio clips are in WAV format, single channel, with a sampling rate of 16 kHz. Sentences have an average duration of 4,815 ms. The intents ’CASH_DEPOSIT’, ’FREEZE’, ’LATEST_TRANSACTIONS’ contain the longest queries, whereas ’AMOUNT’, ’BUS_RESERVATION’, ’OPEN_ACCOUNT’ and ’TRANSFER_MONEY’ generally have shorter query durations. The total duration of the dataset is approximately 4 hours and 17 minutes. Table 3: List of intents. intent domain BALANCE e-banking CASH_DEPOSIT FREEZE LATEST_TRANSACTIONS PAY_BILL OPEN_ACCOUNT TRANSFER_MONEY AMOUNT BUS_RESERVATION transport TECHNICAL_VISIT kontxaalisbëggoonajëflantexam65653737343428282727 Figure 2: Top 5 most frequent words: kont (account), xaalis (money), bëggoona (I wanted to), jëflante (operation), xam (know).
The dataset was collected with the participation of students from Cheikh Anta Diop University in Dakar (UCAD), specifically from the Faculty of Letters and Human Sciences. Each participant recorded their voice using the elicitation mode of the Lig-Aikuma software [Blachon et al., 2016]. A total of 186 utterances were recorded by participants in a controlled environment. The elicitation mode of the Lig-Aikuma software was installed on an Android tablet. At the start of each session, information about the participant are requested, including native language, region of origin (see figure 3), gender, year of birth and name. This information is subsequently stored as metadata on the tablet’s memory card. Note that, the scores presented in figure 3 are not representative of the population of each region. We tried to cover various ethnic groups in the country in order to have diverse accents and dialects as described in Appendix 8.2. Additional details on demographic data and distribution can be found in Appendix 8.1.
To anonymize participant names, the system generates a user_id to replace the actual names (further details on the ethical collection and processing of personal data are provided in Appendices 8.11 and 10). Next, a text file containing the sentences to be pronounced is selected. During recording, sentences were presented one at a time, with the option to cancel if a mispronunciation occurred at any stage. In terms of gender diversity, a total of 31 recording sessions were conducted, including 14 males, 14 females and 3 unspecified. The text and audio data were split with 80% allocated for train set and 20% for test set. Prior to data splitting, preprocessing steps were applied, including the removal of punctuation marks and numbers, and conversion of text to lowercase. Corrupted audio files were also removed from the dataset.
Dakar25.8%Diourbel12.9%Kaolack25.8%Louga6.5%Matam19.4%Thies9.7%
Figure 3: Senegal region repartition by speaker
4 Tasks & Settings
4.1 Automatic Speech Recognition
Automatic Speech Recognition (ASR) refers to the technology that enables a model to recognize and convert spoken language into text. ASR systems are widely used in various applications such as voice assistants, transcription services and customer service automation. Significant research has been conducted in this area, including Listen Attend and Spell by Chan et al. [2016], an end-to-end speech recognition system based on a sequence-to-sequence neural network. Graves et al. [2006] proposed Connectionist Temporal Classification (CTC) model used for training deep neural networks in speech recognition as well as other sequential problems where there is no explicit alignment information between the input and output. More recently, the NVIDIA team Żelasko et al. [2025] developed the Canary Flash model, a variant of Canary models Puvvada et al. [2024] notable for being both multilingual and multitask. This model achieved state-of-the-art results on several benchmarks, including ASR.
Another state-of-the-art model developed by Microsoft and named Phi-4-multimodal-instruct Microsoft et al. [2025] has recently been released to address ASR tasks, as well as vision and text applications. Phi-4-multimodal achieves an average WER score of 6.14% and currently ranks second on Huggingface’s OpenASR leaderboard  [^6] https://huggingface.co/spaces/hf-audio/open_asr_leaderboard. To evaluate ASR models, Word Error Rate (WER) scores are reported in Section 5.4.
4.2 Intent Detection
Intent Detection (ID) is an NLP task that involves classifying a sentence to identify its underlying intent. Several contributions have been made in the field. For instance, Gerz et al. [2021] published MINDS-14, the first training and evaluation dataset for the ID task using spoken data. It includes 14 intents derived from e-banking domain, with spoken examples in 14 different languages. Si et al. [2023a] introduced SpokenWOZ, a large-scale speech-text dataset for spoken Task-Oriented Dialogue in 8 domains and 249 hours of audio. Casanueva et al. [2020] published BANKING77, a dataset comprising 13,083 examples across 77 intents in banking domain.
In this article, we aim to assess the challenging potential of our dataset to better highlight its relevance to the community by evaluating downstream tasks using progressively sophisticated ML Workflows. We first consider classic machine learning algorithms such as k-nearest neighbor (KNN), support vector machine (SVM), linear regression (LR) and naive bayes (NB) as initial baselines. The results are reported in Appendix 8.9. We then consider slightly more sophisticated approaches, using the LASER sentence encoder Heffernan et al. [2022] pre-trained on several languages, including Wolof, for sentence encoding, followed by standard classification models such as Multi-Layer Perception (MLP) and Convolution Neural Networks (CNN). Results are reported in Appendix 8.10. Subsequently, more advanced models based on BERT are evaluated in zero-shot-classification and few-shot-classification mode, and the same models are fine-tuned on the WolBanking77 dataset for comparison. Finally, several benchmarks of recent Small Language Models (SLM) such as Llama-3.2-3B-Instruct  [^7] https://huggingface.co/meta-llama/Llama-3.2-3B-Instruct and Llama-3.2-1B-Instruct  [^8] https://huggingface.co/meta-llama/Llama-3.2-1B-Instruct are presented. To evaluate ID models, F1, precision and recall metrics are reported in Section 5.4.
5 Dataset evaluation & Experiments
In this section, pre-trained models in multiple languages, including African languages, are presented for ID and ASR tasks. Details of the hyperparameter settings are provided in Appendix 8.6. The results demonstrate the value of the WolBanking77 dataset for the community, highlighting its potential to improve ID task in Wolof and to serve as a basis for extending approaches to other low-resource languages.
5.1 Intent detection models
The next sections present and detail the pre-trained models used in the ID task. Experiments for zero and few shot classification are also discussed.
5.1.1 Zero-shot classification with WolBanking77
Zero-shot text classification is a NLP task in which a model, trained on labeled data from specific classes, can classify text from entirely new, unseen classes without additional training. In this experiments WolBanking77 provides the new, unseen dataset. To simulate Zero-shot text classification (more details in Appendix 8.7), following pre-trained models are considered:
BERT base (uncased): A transformers model pre-trained on English introduced by Devlin et al. [2019], using a Masked Language Modeling (MLM) objective.
Afro-xlmr-large: Based on XLM-R-large Conneau et al. [2020] using MLM objective, this model was published by Alabi et al. [2022] and pre-trained on 17 African languages but not Wolof, while still demonstrating in the original paper good scores for NER task on Wolof.
AfroLM_active_learning: Based on XLM-RoBERTa (XLM-R) using MLM objective Ogueji et al. [2021], this model was published by [Dossou et al., 2022]. AfroLM_active_learning is a Self-Active Learning-based Multilingual Pretrained Language Model for 23 African Languages including Wolof.
mDeBERTa-v3-base-mnli-xnli: Based on DeBERTaV3-base He et al. [2023, 2021], this model was published by Moritz et al. [2022] and pre-trained on the CC100 multilingual dataset Conneau et al. [2020], Wenzek et al. [2020] with 100 different languages including Wolof. This model can perform Natural Language Inference (NLI) on 100 languages.
AfriTeVa V2 Base: A multilingual T5 model released by Oladipo et al. [2023] and pre-trained on WURA dataset Oladipo et al. [2023] covering 16 African Languages not including Wolof.
F1-score metrics for Zero-shot text classification are reported in Section 5.4.
5.1.2 Few-shot classification with WolBanking77
Given the results obtained in zero-shot classification, it is relevant to leverage manually annotated data to improve the generalization capacity of existing models on the dataset presented in this article, WolBanking77. All pre-trained models cited in section 5.1.1 are selected for the few-shot classification to measure the gap between large pre-trained models with and without a small number of Wolof samples from WolBanking77 used for fine-tuning. F1-score metrics for few-shot text classification are reported in Section 5.4.
5.2 ASR System
State-of-the-art ASR are selected as baseline models, pre-trained on multiple languages with the possibility of fine-tuning them for a specific language and domain. A description of these models, with additional details, is provided below:
Canary Flash: Canary Flash Żelasko et al. [2025] has been pre-trained on several languages (English, German, French, Spanish) and on various tasks such as ASR and translation. Canary Flash is based on Canary Puvvada et al. [2024] that is an encoder-decoder model whose encoder is based on the FastConformer model Rekesh et al. [2023] and the decoder on the Transformer architecture [Vaswani et al., 2017]. Canary has the distinctive feature of concatenating tokenizers Dhawan et al. [2023] from different languages using SentencePiece. [^9] Google Sentencepiece Tokenizer In this work, English, Spanish, French and Wolof languages are concatenated with Canary’s Tokenizer. These tokens are then transformed into token embedding before being fed into the Transformer decoder. In addition to the tokens of each language, Canary uses 1,152 special tokens representation. At the time of writing, Canary has three variants: canary-1b, canary-1b-flash, and canary-180m-flash. Canary-1b-flash version is chosen for the experiments because of its multilingual support. The canary-1b-flash model was trained on 85K hours of speech data, including 31K hours of public data (FLEURS Conneau et al. [2023], CoVOST v2 Wang et al. [2021b]) and the remainder on private data. The Mozilla CommonVoice 12 dataset Ardila et al. [2020] was used as validation data for each language.
Phi-4-multimodal-instruct: Phi-4-multimodal is a multimodal Small Language Model (SLM) supporting image, text and audio within a single model. It can handle multiple modalities without interference thanks to the Mixture of LoRAs [Hu et al., 2022]. To enable multilingual inputs and outputs, the tiktoken tokenizer  [^10] Tiktoken Tokenizer is used with a vocabulary size of approximately 200K tokens. The model is based on a Transformer decoder Vaswani et al. [2017] and supports a context length of 128K based on LongRopE [Ding et al., 2024]. For the speech/audio modality, several modules have been introduced, including an audio encoder composed of 3 CNN layers and 24 Conformer blocks [Gulati et al., 2020]. To map 1024-dimensional audio features to the 3072-dimensional text embedding space, 2 MLP layers are used in the Audio Projector module. Note that LoRAA was used to all attention and MLP layers with a rank of 320. Phi-4-multimodal was trained on 2M hours of private speech-text pairs in 8 languages. A second post-training phase using Supervised Fine Tuning (SFT) on speech/audio data pairs was then conducted. For the ASR task, this included 20k hours of private data and 20k hours of public data across 8 languages. The SFT data follows the format below:
<|user|><audio>{task prompt}<|end|><|assistant|>{label}<|end|>
The task prompt designates the specific task on which the model is to be fine-tuned.
Phi-4-multimodal outperforms nvidia/canary-1b-flash Żelasko et al. [2025], WhisperV3 1.5B Radford et al. [2023], SeamlessM4T-V2 2.3B Communication et al. [2023], Qwen2-audio 8B Chu et al. [2024], Gemini-2.0-Flash Team et al. [2023] and GPT-4o OpenAI et al. [2024] on the OpenASR dataset leaderboard  [^11] https://huggingface.co/spaces/hf-audio/open_asr_leaderboard with a WER score of 6.14.
Distil-whisper-large-v3.5: Distil-whisper-large-v3.5 is based on Whisper introduced by Radford et al. [2023], which was trained through supervised learning on 680,000 hours of labeled audio data. The authors demonstrated that models trained with this scale could generalize to any dataset through zero-shot learning, meaning they can adapt without requiring fine-tuning for specific datasets. The training data covered 97 different languages. Several versions of Whisper have been released in recent years, including version 3 on which the distil-whisper-large-v3.5 model was built by knowledge-distillation following the methodology described by Gandhi et al. [2023]. Distil-whisper-large-v3.5 was trained on 98k hours of diverse filtered datasets such as Common Voice Ardila et al. [2020], LibriSpeech Panayotov et al. [2015] , VoxPopuli Wang et al. [2021a] , TED-LIUM Hernandez et al. [2018], People’s Speech Galvez et al. [2021], GigaSpeech Chen et al. [2021], AMI Carletta et al. [2006], and Yodas [Li et al., 2023].
5.3 Model training
All models presented in this article are based on Pytorch library [Paszke et al., 2019]. All experiments were conducted on Runpod  [^12] https://www.runpod.io/ and Kaggle. [^13] https://www.kaggle.com/ P100 GPU (16GB VRAM) was used to finetune BERT-Base and mDeBERTa-v3-base. RTX 4090 GPU for Afro-xlmr-large, AfroLM_active_learning, AfriteVa V2 and Llama-3.2 models. For Zero-shot and Few-shot classification, RTX 2000 Ada GPU (16 GB VRAM) was used. Few-shot classification was performed using SetFit huggingface library [Tunstall et al., 2022]. 2-shots refers to 2 samples per intent while 8-shots represents 8 samples per intent. All pre-trained text models are hosted on huggingface platform, they were fine-tuned using transformers library Wolf et al. [2020] except for Llama-3.2 which was fine-tuned using torchtune library [torchtune maintainers, 2024].
MPL and CNN models were trained from scratch for 200 epochs. The MLP model consists of a single 800-dimensional hidden layer, a ReLU activation layer, and a fully connected output layer. For CNN, a Conv1d layer followed by a ReLU activation layer and a fully connected layer for output classification. The Wolbanking77 data was structured in the form of a prompt, as illustrated in Appendix 8.8, to enable the Llama-3.2 model to generate the correct intent.
Canary Flash and Phi-4-multimodal-instruct ASR models were fine-tuned on the Wolbanking77 audio dataset using the NVIDIA NEMO framework Kuchaiev et al. [2019] and Huggingface Transformers library, respectively, on an A100 SXM GPU (80GB VRAM). In contrast, Distil-whisper-large-v3.5 was fine-tuned using the Huggingface Transformers framework on an RTX 2000 Ada GPU (16 GB VRAM). All ASR models were fine-tuned for 1,000 steps.
5.4 Results and Discussions
Figure 4: Zero-shot, few-shot and fine-tuning (FT) results (in % for F1-score) of multilingual pre-trained models on WolBanking77.
BERT-BaseAfroXLMRAfroLMmDeBERTa-v3AfritevaV20505010010011110.20.23311202026261717101014144646474741413535282872727979727269695555F1-scoreScores on 5k data0-shot2-shots8-shotsFT
Figure 4 shows the weighted average F1-score results for zero-shot and few-shot classification. AfroXLMR model outperforms all multilingual pre-trained models in few-shot and fine-tuning settings (with an F1-scores of 79% on 5k samples), followed by BERT-Base and AfroLM. This results also show that existing models, even when pre-trained on Wolof, do not perform well and that our dataset presents specific challenges.
Figure 5: Zero-shot and fine-tuning (FT) results (in %) on multilingual pre-trained models.
BERT-BaseAfroXLMRAfroLMmDeBERTa-v3AfritevaV2050501001002.32.3220.40.41.51.50.50.555555757565652525050F1-scoreScores on all samples0-shotFT
Experiments were conducted on the entire WolBanking77 dataset and reported in figure 5. Similar to the experiments performed on the 5k sub-sample of WolBanking77, results shows that AfroXLMR achieves slightly better scores compared to BERT-Base and AfroLM. These findings highlight that ID task for low-resource languages remains a challenging task, even for state-of-the-art small language models, as shown in Table LABEL:tab:slm.
Table 4: Precision, recall and f1-score scores for Small Language Models on WolBanking77. Best models are highlighted in lightgray.
5k samples All samples
Model Precision Recall F1 Precision Recall F1
Llama-3.2-1B-Instruct 0.74 0.71 0.72 0.52 0.45 0.46
[…]
A comparison was also conducted for ASR models on the audio part of WolBanking77 dataset. Special characters were removed from the text, and all text was converted to lowercase. We can observe from the results with 4 hours of speech data in Table 5 that, Canary-1b-flash with a WER score of 0.59% outperforms Phi-4-multimodal-instruct (WER 3.1%) and more particularly Distil-whisper-large-v3.5 (WER 4.63%). All pre-trained ASR models were fine-tuned on WolBanking77 audio dataset for 1000 steps. The results indicate that strong performance can be achieved with relatively little data, which is promising for WolBanking77 and other low-resource language contexts.
Table 5: Word Error Rates (WER) with 4 hours of speech. Training time is reported in minutes.
Model Parameters Training time Steps WER
Phi-4-multimodal-instruct 5.6B 32 1000 3.1%
Distil-whisper-large-v3.5 756M 44 1000 4.63%
[…]
6 Long-term Support and Future Work
Long term support includes updates to the text and audio recordings, improvements to the simulation setup script, and potential bug fixes. Free access is provided to various versions of the dataset and source code under a CC BY 4.0 license to accelerate research in the field. It is planned to add more audio recordings in diverse environments to enhance model robustness. Text data corresponding to possible responses for each intent will be shared for potential use in Text-To-Speech applications. Additionally, the dataset can be annotated for slot-filling tasks.
7 Conclusion
Access to financial services or public transportation services can be facilitated for illiterate people by providing them with voice interfaces in their language of communication. In this paper, an Intent Detection dataset is presented in Wolof language in two modalities which are voice and text. Additionally, dataset description and benchmarks are presented. WolBanking77 is published and shared freely under CC BY 4.0 license along with the code and datasheets Gebru et al. [2021] with the aim of inspiring low budget research into low-resource languages.
Acknowledgments and Disclosure of Funding
This work is supported by the Partnership for Skills in Applied Sciences, Engineering and Technology (PASET) - Regional Scholarship and Innovation Fund (RSIF) under Grant No.:B8501L10014. We also thank the entire CLAD team and in particular: Professor Souleymane Faye and all the translators who participated in this project.